2tail test with rejecting region Alfa_.050, n=16 d.f.15, sample mean .950, standar deviation =.01
Ho' m=1
Hi m =/ 1

Answer by stanbon(75887)

2tail test with rejecting region Alfa_.050, n=16 d.f.15, sample mean .950, standard deviation =.01
-----
Ho' m=1
Hi m =/ 1
------
t(0.95) = (0.95-1)/[0.01/sqrt(16)] = -20
----
p-value = 2*P(t < -20 when df = 15) = 2*tcdf(-100,-20,15) = 3.16x10^-12
=============================
Conclusion: Reject Ho at the 5% level.
=============================
Cheers,
Stan H.